# Imitation from Observations with Trajectory-Level Generative Embeddings

### v1

@@ Abstract
We consider the offline imitation learning from observations (LfO) where the expert demonstrations are scarce and the available offline suboptimal data are far from the expert behavior. Many existing distribution-matching approaches struggle in this regime because they impose strict support constraints and rely on brittle one-step models, making it hard to extract useful signal from imperfect data. To tackle this challenge, we propose TGE, a trajectory-level generative embedding for offline LfO that constructs a dense, smooth surrogate reward by estimating expert state density in the latent space of a temporal diffusion model trained on offline trajectory data. By leveraging the smooth geometry of the learned diffusion embedding, TGE captures long-horizon temporal dynamics and effectively bridges the gap between disjoint supports, ensuring a robust learning signal even when offline data is distributionally distinct from the expert. Empirically, the proposed approach consistently matches or outperforms prior offline LfO methods across a range of D4RL locomotion and manipulation benchmarks.
@@ 1 Introduction
Imitation learning (IL) has demonstrated remarkable success in enabling agents to acquire complex behaviors across various domains, from robotic manipulation (florence2022implicit; chi2023diffusion; brohan2023rt1) to autonomous navigation (codevilla2018end; bansal2019chauffeurnet). By directly supervising state-action mappings, standard IL methods can effectively recover high-performance policies. However, applying these techniques to real-world scenarios poses substantial challenges. A primary bottleneck is the absence of explicit action information in many natural data sources, such as video demonstrations or third-person recordings, which precludes direct supervision of state-action mappings (torabi2018behavioral; stadie2017third; liu2018imitation). This has led to the emergence of Learning from Observations (LfO), where the agent must infer desirable behaviors from state-only expert trajectories.
@@
While some LfO methods attempt to recover latent actions via inverse dynamics or auxiliary inference mechanisms (hanna2017grounded; kidambi2020morel), a second, equally critical challenge remains: the quality of available data. In offline settings, the large datasets required for training often contain suboptimal or noisy behaviors arising from human inconsistency, task ambiguity, or domain shifts (levine2020offline; mandlekar2022matters). Effectively exploiting such imperfect data requires algorithms that can filter noise and extract valuable behavioral priors without being misled by non-expert trajectories.
@@
Despite steady progress, current Learning from Observations (LfO) methods confront structural weaknesses that limit their robustness. Surrogate reward approaches, such as ORIL (zolna2020offline), rely on discriminator or inverse dynamics models to annotate offline data, but their performance can deteriorate over long horizons when the learned signal is imperfect (ross2011reduction). On the other hand, state-of-the-art distribution-matching approaches typically rely on estimating density ratios between expert and offline distributions (ho2016generative; ma2022versatile). However, these methods typically assume the expert behavior is well-covered by the offline data. When distributions are disjoint, strictly aligning them yields sparse or undefined signals, failing to guide the agent towards the expert region. We review these paradigms in more detail in Sec. 2.1.
@@
To address these limitations, we propose Trajectory-level Generative Embeddings (TGE), a novel framework for offline LfO that bypasses the structural reliance on support coverage inherent in density ratio estimation. Instead of relying on brittle density ratios, TGE leverages the smooth, low-dimensional manifold learned by a trajectory-based diffusion model (janner2022planning). By encoding trajectories into this generative latent space, we construct a dense surrogate reward signal that bridges disjoint supports. We specifically employ a Laplacian kernel with heavy-tailed decay for reward estimation, which ensures an informative reward signal throughout the suboptimal data, preventing the reward collapse issues common in prior distance-based methods. This reward signal effectively guides offline RL algorithms, such as IQL (kostrikov2021offline) or ReBRAC (tarasov2023revisiting), toward the expert policy.
@@
  • We introduce a trajectory-level representation learning approach based on a generative planner. We find that the resulting latent embeddings are naturally separable and can distinguish between suboptimal and expert-like data in offline datasets, without requiring explicit discriminative training.
@@ 2.1 Offline Imitation Learning
Offline imitation learning (IL) aims to recover an expert policy from a fixed dataset of demonstrations without access to online environment interactions. While the foundational approach, Behavioral Cloning (BC) (pomerleau1989alvinn), treats this as a supervised regression problem, it notoriously suffers from distribution shift and compounding errors when the agent accesses out-of-distribution states (ross2011reduction). To mitigate these compounding errors, distribution-matching approaches shift the objective from local action supervision to global distribution alignment, minimizing the divergence between the learner and expert occupancy measures (kostrikov2020imitation; garg2021iq). Recent advances have largely shifted toward conditional generative modeling to better capture the multimodal nature of the expert’s behavior. Prominent methods now leverage Energy-Based Models (florence2022implicit), Transformers (shafiullah2022behavior), and Diffusion Models (chi2023diffusion) to represent complex, non-Gaussian action distributions, thereby enabling robust policy synthesis even in high-dimensional and multimodal settings.
@@
Learning from Observations (LfO) arises as a critical and more challenging subset of this paradigm, where the expert demonstrations lack action labels. Offline LfO approaches generally fall into two distinct paradigms based on how they utilize expert data:
Occupancy Matching
@@
This class of methods formulates imitation as a convex optimization problem, solving for the state-occupancy ratio $w(s)=d^{\pi_{E}}(s)/d^{\pi}(s)$ between the expert and offline data distributions. Methods like SMODICE (ma2022versatile) and LobsDICE (kim2022lobsdice) minimizing the KL-divergence between these distributions are often sensitive to the distribution of the offline buffer. Recent variants like PW-DICE (yan2024offline) introduces Wasserstein geometry to relax the strict support constraints, and DILO (sikchi2024dual) utilizes a dual formulation of the $\chi^{2}$-divergence to improve stability. Despite these advances, these methods fundamentally rely on estimating density ratios, which can become ill-conditioned when the offline and expert supports are disjoint.
Surrogate Reward Learning
@@
Alternatively, two-stage approaches first learn a surrogate reward function $r(s)$ from data and then optimize a policy using offline RL. This paradigm extends beyond imitation learning. It is central to unsupervised reinforcement learning (URLB) (laskin2021urlb), where methods typically derive intrinsic pseudo-rewards, such as those based on particle entropy or diversity, to guide policy acquisition in the absence of external supervision. In the context of offline LfO, standard approaches like ORIL (zolna2020offline) employ adversarial learning to train a discriminator-based reward. However, adversarial objectives are notoriously unstable and often produce signals that are difficult to optimize. TGE falls into this surrogate reward paradigm but circumvents these stability issues by replacing adversarial discriminators with a dense, geometric distance metric derived from latent space of a trajectory-level generative planner trained on the offline dataset.
@@ 2.2 Diffusion Models in Reinforcement Learning
Diffusion models have rapidly been adopted in RL, primarily for generative control and data synthesis. In the control domain, diffusion models function as policies (chi2024diffusionpolicy; wangdiffusion) or trajectory planners (janner2022planning; ajay2022conditional), leveraging iterative denoising to capture multimodal distributions and solve long-horizon tasks. Recent works have extended this to value-based methods, using diffusion to regularize offline Q-learning (hansen2023idql). Beyond control, diffusion models are increasingly used as world models for hallucinating environments (alonso2024diffusion). Our work focuses on a distinct application: metric learning. While methods like Stable Rep (tian2023stablerep) or Diffusion Reward (huang2024diffusion; wang2025goal) similarly derive learning signals from generative models, they typically rely on static visual semantics from video or image models. In contrast, TGE exploits the trajectory-level latent embedding of a temporal diffusion encoder to capture system dynamics, augmenting the reward-free offline dataset that enables robust imitation without the computational cost of generative sampling during downstream offline RL training.
@@ 2.3 Representation Learning in Reinforcement Learning
Representation learning is central to scaling reinforcement learning to high-dimensional observation spaces by extracting compact embeddings that filter task-irrelevant noise. Early approaches largely relied on reconstruction-based auxiliary tasks, utilizing world models or autoencoders to compress sensory inputs into latent states that capture the factors of the environment ha2018world. To improve robustness to task-irrelevant visual details, subsequent research adopted contrastive learning and bisimulation metrics, which optimize for invariance by grouping states with similar transition dynamics laskin2020curl; zhang2020learning. However, these approaches typically operate on static state snapshots, which implicitly assumes that single-frame observations are sufficient to distinguish system states. To address this, we depart from single-step encoders and instead employ diffuser janner2022planning as an encoder to capture trajectory-level representations to provide the temporal context necessary to distinguish behaviors that appear visually identical but possess divergent underlying dynamics.
@@ 3 Preliminaries
We formulate the learning problem within the standard Markov Decision Process (MDP) framework, defined by the tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},P,r,\gamma\rangle$, where $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, respectively. The transition kernel $P(s^{\prime}|s,a)$ specifies the probability of transitioning to a next state $s^{\prime}$ given the current state $s$ and action $a$. The reward function $r(s,a)$ provides scalar feedback for each state–action pair, and $\gamma\in[0,1)$ is a discount factor that determines the weighting of future rewards. A policy $\pi(a|s)$ defines a distribution over actions conditioned on the state, and the objective of reinforcement learning is to identify a policy that maximizes the expected discounted return $J(\pi)=\mathbb{E}_{\pi,P}\!\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right].$ In imitation learning or learning from observations, both the true reward function $r$ and the expert actions are typically unavailable. Consequently, the learner must infer desirable behaviors from state-only expert trajectories and suboptimal data providing limited state–action coverage.
@@
Figure 1: Overview of Trajectory-level Generative Embeddings (TGE). The framework employs a trajectory-level diffusion encoder to map trajectory segments into a latent embedding space. A surrogate reward is then computed using a Laplacian kernel over these embeddings to augment the reward-free suboptimal dataset, enabling offline RL training for policy learning.
Offline Imitation Learning from Observations
@@
We consider the learning from observation (LfO) setting, where the small expert dataset contains only state observations, without any expert actions or reward annotations. In addition, our work focuses on the offline imitation learning regime, where the agent cannot directly interact with the environment. Instead, we assume access to an offline dataset composed of suboptimal state–action trajectories, which provides the necessary coverage for learning.
@@
Let the expert dataset containing a single expert trajectory be denoted by $\mathcal{D}_{E}$, and let $\mathcal{D}_{\mu}$ denote the suboptimal offline dataset consisting of state–action transitions generated by a mixed behavioral policy $\mu$, which contains a combination of expert and suboptimal trajectories. The agent does not have access to the trajectory labels; that is, it cannot directly distinguish which trajectories are expert and which are suboptimal. This setting is standard practice in offline imitation learning from observations (sikchi2024dual; yan2024offline). We define the normalized discounted state visitation distribution of a policy $\pi$ as $\rho_{\pi}(s)=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}P(s_{t}=s|\pi)$. Accordingly, let $\rho_{E}$, $\rho_{\pi}$, and $\rho_{\mu}$ represent the distributions induced by the expert policy $\pi_{E}$, the learned policy $\pi$, and the data-collecting policy $\mu$, respectively. The goal of the learning algorithm is to recover a policy $\pi$ whose behavior closely matches that of the expert policy $\pi_{E}$.
Denoising Diffusion Probabilistic Models
@@
Denoising Diffusion Probabilistic Models (DDPMs) (ho2020denoising) are a class of generative models that learn to approximate complex data distributions through a sequence of iterative denoising operations. A DDPM defines a forward diffusion process that gradually corrupts a clean sample $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$ by adding Gaussian noise over $T$ time steps according to $q(\mathbf{x}_{k}\mid\mathbf{x}_{k-1})=\mathcal{N}\!\left(\sqrt{1-\beta_{k}}\,\mathbf{x}_{\,k-1},\,\beta_{k}I\right)$, where $\{\beta_{k}\}_{k=1}^{K}$ is a variance schedule controlling the noise magnitude at each step. After sufficient diffusion steps, this process yields a nearly isotropic Gaussian distribution $q(\mathbf{x}_{K})\approx\mathcal{N}(0,I)$.
@@
The generative (reverse) process is parameterized by a neural network $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{k},k)$ that predicts the injected noise and induces the reverse transitions $p_{\theta}(\mathbf{x}_{k-1}\mid\mathbf{x}_{k})=\mathcal{N}\!\left(\mu_{\theta}(\mathbf{x}_{k},k),\,\Sigma_{\theta}(\mathbf{x}_{k},k)\right)$, where the mean $\mu_{\theta}(\mathbf{x}_{k},k)$ is computed from the noise prediction $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{k},k)$. Training is performed by minimizing a reweighted variational lower bound, which simplifies to the following mean-squared error objective:
@@
$\mathcal{L}=\mathbb{E}_{k,\,\mathbf{x}_{0},\,\boldsymbol{\epsilon}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\!\left(\sqrt{\bar{\alpha}_{k}}\,\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{k}}\,\boldsymbol{\epsilon},\,k\right)\right\|_{2}^{2}\right],$
@@
where $\bar{\alpha}_{k}=\prod_{i=1}^{k}(1-\beta_{i})$. Once trained, the model generates new samples by starting from Gaussian noise $\mathbf{x}_{K}\sim\mathcal{N}(0,I)$ and iteratively denoising through the learned reverse process. Recent advances in offline reinforcement learning have leveraged diffusion models as trajectory-level generative planners, enabling long-horizon reasoning and multi-modal trajectory generation (janner2022planning; ajay2022conditional).

### v2

@@ Abstract
We consider the offline imitation learning from observations (LfO), where expert demonstrations are scarce and contain only state observations, and the suboptimal policy is far from expert behavior. In this regime, many existing imitation learning approaches struggle to extract useful information from imperfect data since they impose strict support constraints and rely on brittle one-step models. To tackle this challenge, we propose Trajectory-level Generative Embedding (TGE) for offline LfO. TGE constructs a dense, smooth surrogate reward by using particle based entropy estimation to maximize the log-likelihood of expert trajectories in the latent space of a temporal diffusion model trained on offline suboptimal data. By leveraging the structured geometry of the learned diffusion embedding, TGE captures long-horizon temporal dynamics and effectively bridges the gap under severe support mismatch, ensuring a robust learning signal even when offline data is distributionally distinct from the expert. Empirically, the proposed approach consistently matches or outperforms prior offline LfO methods across a range of D4RL locomotion and manipulation benchmarks.
@@ 1 Introduction
Imitation learning (IL) has demonstrated remarkable success in enabling agents to acquire complex behaviors across various domains, from robotic manipulation (Florence et al., 2021; Chi et al., 2023; Brohan et al., 2023) to autonomous navigation (Codevilla et al., 2018; Bansal et al., 2019). While standard IL methods can effectively recover high-performance policies by supervising explicit action labels, applying these techniques to real-world scenarios poses substantial challenges. A primary bottleneck is the absence of explicit action information in many natural data sources, such as video demonstrations or third-person recordings, which precludes direct supervision of state-action mappings (Torabi et al., 2018; Stadie et al., 2017; Liu et al., 2018). This has led to the emergence of Learning from Observations (LfO), where the agent must infer desirable behaviors from state-only expert trajectories.
@@
To recover the underlying actions, many of the existing LfO methods attempt to estimate the inverse dynamics or auxiliary inference mechanisms (Hanna and Stone, 2017; Kidambi et al., 2020). However, an equally critical challenge lies in the quality of the offline data itself. As large-scale offline datasets inevitably contain suboptimal or inconsistent behaviors arising from human error or task ambiguity, directly applying behavioral cloning to such datasets often leads to poor performance. Effectively leveraging these imperfect demonstrations therefore requires algorithms that can robustly filter noise and distill informative behavioral priors, without being misled by non-expert trajectories.
@@
In addition to the challenge in the offline data quality, current LfO approaches confront structural weaknesses that limit their robustness despite their progress. To mention a few, surrogate reward approaches, such as ORIL (Zolna et al., 2020), rely on discriminator or inverse dynamics models to annotate offline data. However, their performance can deteriorate over long horizons when the learned signal is imperfect (Ross et al., 2011). On the other hand, state-of-the-art distribution-matching approaches rely on estimating density ratios between expert and offline distributions (Ho and Ermon, 2016; Ma et al., 2022). However, these methods typically assume the expert policy is well-covered by the offline data. When the offline behavioral data have limited support to the expert, strictly aligning them yields sparse or undefined signals and therefore fails to guide the agent towards the expert behavior. We review these paradigms in more detail in Sec. 2.1. These limitations raise a central question for learning from observations:
How can we extract a reliable imitation signal from imperfect offline data under support mismatch?
@@
To address these limitations, we propose Trajectory-level Generative Embeddings (TGE), a novel framework for offline LfO that bypasses the structural reliance on support coverage inherent in density ratio estimation. Instead of relying on unstable density ratios, TGE leverages the smooth representation learned by a trajectory-based diffusion model such as diffuser (Janner et al., 2022). By encoding trajectories into this generative latent space, we construct a dense surrogate reward signal that bridges regions of sparse support. Specifically, we employ a logarithmic distance kernel with heavy-tailed decay for reward estimation inspired by the particle-based entropy estimation Liu and Abbeel (2021); Mutti et al. (2020); Singh et al. (2003), which ensures that maximizing the surrogate reward using standard offline RL algorithms, such as IQL (Kostrikov et al., 2021) or ReBRAC (Tarasov et al., 2023) is essentially maximizing the log-likelihood of the expert policy. Notably, this design of the surrogate reward also ensures an informative signal throughout the suboptimal data, preventing the reward collapse issues common in prior distance-based methods.
@@
  • We introduce a trajectory-level representation learning approach based on a generative planner. We demonstrate that the resulting latent embeddings are naturally separable and can distinguish between suboptimal and expert-like data in offline datasets, without requiring explicit discriminative training.
@@ 2.1 Offline Imitation Learning
Offline imitation learning (IL) aims to recover an expert policy from a fixed dataset of demonstrations without access to online environment interactions. While the foundational approach, Behavioral Cloning (BC) (Pomerleau, 1989), treats this as a supervised regression problem, it notoriously suffers from distribution shift and compounding errors when the agent accesses out-of-distribution states (Ross et al., 2011). To mitigate these compounding errors, distribution-matching approaches shift the objective from local action supervision to global distribution alignment, minimizing the divergence between the learner and expert occupancy measures (Kostrikov et al., 2020; Garg et al., 2021). Beyond improving the training objective, recent advances have emphasized policy expressivity, shifting toward conditional generative modeling to better capture the multi-modal nature of the expert’s behavior. Prominent methods now leverage Energy-Based Models (Florence et al., 2021), Transformers (Shafiullah et al., 2022), and Diffusion Models (Chi et al., 2023) to represent complex, non-Gaussian action distributions, thereby enabling robust policy synthesis even in high-dimensional and multi-modal settings.
@@
Learning from Observations (LfO) arises as a critical and more challenging subset of this paradigm, where the expert demonstrations lack action labels. Offline LfO approaches generally fall into two distinct paradigms based on how they utilize expert data which we provide an overview below:
@@
Occupancy Matching. This class of methods formulates imitation as a optimization problem, solving for the state-occupancy ratio $w(s)=d^{\pi_{E}}(s)/d^{\pi}(s)$ between the expert and offline suboptimal data distributions. Methods like SMODICE (Ma et al., 2022) and LobsDICE (Kim et al., 2022), which minimize the KL-divergence between these distributions, are often sensitive to the distribution of the offline buffer. Recent variants like PW-DICE (Yan et al., 2024) introduce Wasserstein geometry to relax the strict support constraints, while DILO (Sikchi et al., 2025) utilizes a dual formulation of the $\chi^{2}$-divergence to improve stability. Despite these advances, these methods fundamentally rely on estimating density ratios, which can become ill-conditioned when the offline behavioral data has limited support to the expert policy.
@@
Surrogate Reward Learning. Another line of work is to introduce the surrogate rewards guiding the agent to align with the expert policy. Usually it contains a two-stage approach in which the agent first learns a surrogate reward function $r(s)$ from data and then optimizes a policy using offline RL algorithms. Broadly speaking, this paradigm extends beyond imitation learning and is central to unsupervised reinforcement learning (URLB) (Laskin et al., 2021), where methods typically derive intrinsic pseudo-rewards to guide policy acquisition or environment exploration in the absence of external supervision. In the context of offline LfO, standard approaches like ORIL (Zolna et al., 2020) apply adversarial learning to train a discriminator-based reward. However, adversarial objectives are notoriously unstable and often produce signals that are difficult to optimize. Remarkably, our method TGE falls into this surrogate reward paradigm but circumvents these stability issues by replacing adversarial discriminators with a dense, geometric distance metric derived from the latent space of a trajectory-level generative planner trained on the offline dataset.
@@ 2.2 Diffusion Models in Reinforcement Learning
Diffusion models have rapidly been adopted in RL, primarily for generative control and data synthesis. In the control domain, diffusion models function as policies (Chi et al., 2025; Wang et al., ; Ding et al., 2024) or trajectory planners (Janner et al., 2022; Ajay et al., 2022), leveraging iterative denoising to capture multi-modal distributions and solve long-horizon tasks. Recent works have extended this to value-based methods, using diffusion to regularize offline Q-learning (Hansen-Estruch et al., 2023). Beyond control, diffusion models are increasingly used as world models for hallucinating environments (Alonso et al., 2024). In the context of diffusion models in RL, our work focuses on a distinct application in using diffusion model for metric learning. While methods like Stable Rep (Tian et al., 2023) or Diffusion Reward (Huang et al., 2024) similarly derive learning signals from generative models, they typically rely on static visual semantics from video or image models. In contrast, TGE exploits the trajectory-level latent embedding of a temporal diffusion encoder to capture system dynamics, augmenting the reward-free offline dataset that enables robust imitation without the computational cost of generative sampling during downstream offline RL training.
@@ 2.3 Representation Learning in Reinforcement Learning
Representation learning is central to scaling RL to high-dimensional observation spaces by extracting compact embeddings that filter task-irrelevant noise. Early approaches largely relied on reconstruction-based auxiliary tasks with world models or autoencoders to compress sensory inputs into latent states that capture the factors of the environment Ha and Schmidhuber (2018). Subsequent research adopted contrastive learning and bi-simulation metrics to group states with similar transition dynamics Laskin et al. (2020); Zhang et al. (2021) to improve the model robustness to task-irrelevant details. However, these approaches typically learn the representation for single-stage observation. To tackle this, we instead leverage the representation learnt from a trajectory diffusion model (i.e., Diffuser Janner et al. (2022)). These trajectory-level representations provide the temporal context necessary to distinguish behaviors that are visually similar but possess divergent underlying dynamics.
In supporting this line, recent works have shown that diffusion models can be viewed as denoising autoencoders where intermediate features serve as robust representations for downstream tasks (Fuest et al., 2024). For instance, RepFusion (Yang and Wang, 2023) shows that distilling diffusion model yields high-quality representations on image recognition. This suggests that the diffusion training induces semantically meaningful features for complex dynamics.
@@ 3 Preliminaries
Markov Decision Process. We consider the discounted Markov Decision Process (MDP) setting, defined by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, respectively. The transition kernel $P(s^{\prime}|s,a)$ specifies the probability of transitioning to a next state $s^{\prime}$ given the current state $s$ and action $a$ with the reward function defined by $r(s,a)$ and $\gamma\in[0,1)$ be a discount factor. A policy $\pi(a|s)$ defines a distribution over actions conditioned on the state, and the objective of reinforcement learning is to identify a policy that maximizes the expected discounted return $J(\pi)=\mathbb{E}_{\pi,P}\!\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right].$
In the imitation learning from observations setting, both the true reward function $r$ and the expert actions are typically unavailable. Consequently, the learner must infer desirable behaviors from state-only expert trajectories and suboptimal data providing limited state–action coverage.
@@
Figure 1: Overview of Trajectory-level Generative Embeddings (TGE). The framework employs a trajectory-level diffusion encoder to map trajectory segments into a latent embedding space. A surrogate reward is then computed using a logarithmic kernel over these embeddings to augment the reward-free suboptimal dataset, enabling offline RL training for policy learning.
@@
Offline Imitation Learning from Observations. In imitation learning from observation (LfO), the agent is given a small expert dataset containing only state observations, without any expert actions or reward annotations. In addition, our work focuses on the offline learning, where the agent cannot directly interact with the environment. Instead, we assume the agent can access to an offline dataset composed of suboptimal state–action trajectories, which provides the necessary coverage for learning.
@@
Let the expert dataset containing one complete state-only expert episode be denoted by $\mathcal{D}_{E}$, and let $\mathcal{D}_{\mu}$ denote the suboptimal offline dataset consisting of state–action trajectories generated by a mixed behavioral policy $\mu$, where $\mathcal{D}_{\mu}$ contains a combination of expert and suboptimal trajectories. The agent does not have access to the trajectory labels; that is, it cannot directly distinguish which trajectories are expert and which are suboptimal. This setting is standard practice in offline imitation learning from observations (Sikchi et al., 2025; Yan et al., 2024). We define the normalized discounted state visitation distribution of a policy $\pi$ as $\rho_{\pi}(s)=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}P(s_{t}=s|\pi)$. Accordingly, let $\rho_{E}$, $\rho_{\pi}$, and $\rho_{\mu}$ represent the distributions induced by the expert policy $\pi_{E}$, the learned policy $\pi$, and the data-collecting policy $\mu$, respectively. The goal of the learning algorithm is to recover a policy $\pi$ whose behavior closely matches that of the expert policy $\pi_{E}$.
@@
Denoising Diffusion Probabilistic Models. Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) are a class of generative models that learn to approximate complex data distributions through a sequence of iterative denoising operations. In particular, the forward diffusion process gradually corrupts a clean sample $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$ by adding Gaussian noise over $K$ steps according to $q(\mathbf{x}_{k}\mid\mathbf{x}_{k-1})=\mathcal{N}\!\left(\sqrt{1-\beta_{k}}\,\mathbf{x}_{\,k-1},\,\beta_{k}I\right)$, where $\{\beta_{k}\}_{k=1}^{K}$ is a variance schedule controlling the noise magnitude at each step. After sufficient diffusion steps, this process yields a nearly isotropic Gaussian distribution $q(\mathbf{x}_{K})\approx\mathcal{N}(0,I)$.
@@
The generative (reverse) process is parameterized by a neural network $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{k},k)$ that predicts the injected noise and the posterior $p_{\theta}(\mathbf{x}_{k-1}|\mathbf{x}_{k})=\mathcal{N}(\boldsymbol{\mu}_{\theta}(\mathbf{x}_{k},k),\boldsymbol{\Sigma}_{\theta}(\mathbf{x}_{k},k))$, where the mean $\boldsymbol{\mu}_{\theta}(\mathbf{x}_{k},k)$ is computed from the denoising model $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{k},k)$. Training is performed by minimizing a reweighted variational lower bound, which simplifies to the following mean-squared error objective:
@@
$\mathcal{L}=\mathbb{E}_{k,\mathbf{x}_{0},\boldsymbol{\epsilon}}\left[\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{k}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{k}}\boldsymbol{\epsilon},k)\|_{2}^{2}\right],$
@@
where $\bar{\alpha}_{k}=\prod_{i=1}^{k}(1-\beta_{i})$. Once trained, the model generates new samples by iteratively denoising through the learned reverse process starting from Gaussian noise $\mathbf{x}_{K}\sim\mathcal{N}(0,I)$. Recent advances in offline reinforcement learning have leveraged diffusion models as trajectory-level generative planners, enabling long-horizon reasoning and trajectory generation (Janner et al., 2022; Ajay et al., 2022).